OpenStudy (anonymous):

Please help me solve this: I need to find the quadratic equation of a parabola with a focus of (1, 3) and a directrix of y = -3.

OpenStudy (cybershadow):

eh i thought we solved it, lol , wasnt dat correct :S

OpenStudy (anonymous):

Please, someone help me. I've been on this test for an hour and I'm about to cry. x_x

OpenStudy (anonymous):

say (x, y) is some point on the parabola

OpenStudy (anonymous):

@CyberShadow that was a different question. :/

OpenStudy (anonymous):

by definition, distance between any point on parabola and Focus = distance between Focus and directrix

OpenStudy (anonymous):

@beccaboo022a Is there a formula I can use to input values? I know there's one that's something like y = (___)^2 + __

OpenStudy (anonymous):

Would the point between the focus and directrix be the vertex?

OpenStudy (anonymous):

yes there is a formula if u prefer formula to actually solving

OpenStudy (anonymous):

Look, here's the actual question:

OpenStudy (anonymous):

I can use the y = formula right? I jsut can't find the name of that formula and I don't remember what it was called. -____-

OpenStudy (cybershadow):

y = 4ax

OpenStudy (cybershadow):

*y^2

OpenStudy (cybershadow):

if dats the one u r talking abt? :S

OpenStudy (anonymous):

an example equation would be y = 1/4(x-1)^2 + 2. What is the name of this formula and how can I input the directrix and focus into that to get it like that?

OpenStudy (anonymous):

I have to leave in 30 minutes and I've already been stuck on this question for 30. :/

OpenStudy (cybershadow):

see we have to find vertex first that wud be midway of directrix and focus

OpenStudy (anonymous):

Would the vertex be (1, 0)?

OpenStudy (cybershadow):

yea

OpenStudy (cybershadow):

now just use this equation of parabola

OpenStudy (anonymous):

Can you tell me the name of this formula? y = 1/4(x-1)^2 + 2 so i can find the real formula with the letters or w/e

OpenStudy (anonymous):

Wait, I found it. y = a(x - h)^2 + k

OpenStudy (cybershadow):

its (y-h) = 4a(x-k)^2

OpenStudy (cybershadow):

yes!!

OpenStudy (cybershadow):

so now its solved!

OpenStudy (anonymous):

Are they asking for a specific formula in the question? i can't afford to get it wrong :/

OpenStudy (anonymous):

nope, they are just asking u find the equation of quadratic

OpenStudy (anonymous):

Okay, so I can use the y = a(x - h)^2 + k?

OpenStudy (cybershadow):

formula is basically this -(y-h) = 4a(x-k)^2 equation of parabola

OpenStudy (anonymous):

First figure out which direction the quadratic faces ?

OpenStudy (cybershadow):

then u can solve it and rearrange it

OpenStudy (anonymous):

left/right or up/down ?

OpenStudy (anonymous):

How do I find that out? And why can't I just use y = a(x - h)^2 + k? D:

OpenStudy (anonymous):

once u figure that out, look at the attached chart

OpenStudy (anonymous):

I don't know how to find a..

OpenStudy (anonymous):

to figure out which direction it faces : look at the directrix

OpenStudy (anonymous):

@beccaboo022a you're moving way too fast for me. :/

OpenStudy (cybershadow):

a is distance between directrix and vertex

OpenStudy (anonymous):

that's 6 points

OpenStudy (anonymous):



OpenStudy (anonymous):

The focus it ouside and the directrix is in, right?

OpenStudy (cybershadow):

its right handed parabola as vertex is (1,0)

OpenStudy (anonymous):

Alright ! draw the directrix and focus first

OpenStudy (anonymous):

Wait, I'm cmixed up. The directrix is outside and the focus is inside.

OpenStudy (anonymous):

draw it out

OpenStudy (cybershadow):

yeah rite Jay

OpenStudy (anonymous):

Which means that the parabola looks like a smiley face.

OpenStudy (anonymous):

draw it out

OpenStudy (anonymous):